40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4608     Entry time: Tue May 3 10:41:35 2011
Author: josephb 
Type: Update 
Category: CDS 
Subject: Morning maintenance 

1) Filled in the C1SUS_BS_OLMATRIX properly so as to make the BS oplev work for Steve.

2) Turned on the ITMX damping.  Apparently it had tripped this morning, possibly due to work in the lab area.

3) The ETMX FE controller (c1scx) had ADC timed out and died sometime around 8:30 am.  The c1x01 (the IOP on the ETMX computer) was also indicating a FB status error (mismatch in DAQ channels).

The reported error in dmesg on c1iscex was:

[1628690.250002] c1spx: ADC TIMEOUT 0 3541 21 3605
[1628690.250002] c1scx: ADC TIMEOUT 0 3541 21 3605

Just to be safe, I rebuilt the c1x01 and c1scx models, ran ./activateDAQ.py, and used the scripts killc1spx, killc1scx, and killc1x01.

I finally restarted the process with startc1x01, startc1scx, and startc1spx.  Everything is currently alive and indicating all green.

ELOG V3.1.3-